[Hallc_running] [New Logentry] RC status

mack at jlab.org mack at jlab.org
Sat Mar 23 09:55:02 EDT 2019

Logentry Text:
Current Status
Beam checkout plan, superharp scans, and energy measurement done.  (Experts are invited to comment.) As for calibration runs: 1) sieve runs complete, 2) Lumi scans were truncated due to inability to get higher beam currents on Owl, and 3)Hydrogen elastic singles are IN PROGRESS.  Since about 7am, the MCC has been able to deliver 70 muA as requested. 

Plans - Finish calibration runs in sections 3 and start those in section 4 (elastic singles and elastic coincidence, respectively). Need feedback from the spokes on priority for going back and finishing the lumi scan (in the next few hours before we reverse polarity for the SHMS). Shift leaders keep track of the requested mix of LH2 and 10cm Dummy runs, and pass your completion status down to the next shift. 

Announcement - No RC meeting today. I'll chat with the spokes later. 

Kaon LT run plans : https://hallcweb.jlab.org/wiki/index.php/Main_Page#Kaon_LT_experiment
Information for Shift Takers: https://hallcweb.jlab.org/wiki/index.php/Information_for_Shift_Takers


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3671371
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190323/1e1f5e5d/attachment.html>

More information about the Hallc_running mailing list